OpenStudy (anonymous):

if f(x)=x/x^2-16, what is the value of f(-10)

OpenStudy (anonymous):

f(x)=x/x^2-16 f(-10)=-10/(-10)^2-16 =-10/100-16 =-1/10-16 =(-1-16*10)/10 =-161/10 =-16.1

OpenStudy (anonymous):

u didn't mean f(x)=x/(x^2-16) did u

OpenStudy (anonymous):

yes,i meant the second way

OpenStudy (anonymous):

ok f(x)=x/(x^2-16) f(-10)=-10((-10)^2-16)=-10(100-16)=-10/84=-5/42=-0.11

OpenStudy (anonymous):

ok, thanks!